# thanks: Corresponding author

### v1

@@ II.4 Cholesky Decomposition
In the CD framework[47], the two-electron repulsion integrals (ERI) can be efficiently approximated as,
@@
where $\mu$, $\nu$, $\kappa$, $\lambda$ correspond to atomic spinor indices, $L_{\mu\nu}^{P}$ denotes the Cholesky vectors and $n_{CD}$ specifies their dimensionality. For the generation of the Cholesky vector, we employed a one-step algorithm where the CD are constructed iteratively by selecting the largest diagonal element of the ERI matrix $(\mu\nu\mid\mu\nu)$, and the decomposition continues until the maximum remaining diagonal elements falls below a predefined CD threshold[48]. The Cholesky vectors in the AO basis can be transformed to the MO basis as follows,
@@
These transformed vectors can then be used to generate anti-symmetrized two-electron integrals on the fly in the molecular orbital (MO) basis.
@@ II.5 Frozen Natural Spinors (FNS)
The lack of spin symmetry and the need to store complex numbers make relativistic coupled cluster calculations significantly more costly compared to their non-relativistic counterparts. Moreover, one often needs to uncontract the basis set to fit the small component of the Hamiltonian in relativistic calculations, which further increases the cost of relativistic coupled cluster calculations. One of the most effective ways to reduce the computational cost of relativistic coupled cluster calculations is to truncate the virtual space for the calculation using natural spinors[49, 50]. The natural spinors are the relativistic analogs of the natural orbitals introduced by Löwdin[51] and can be obtained by diagonalizing the one-body correlated reduced density matrix (1-RDM), which is constructed from a spin-orbit coupled wave function calculated through some approximate electron correlation calculations[52]. Among the various flavors of natural orbital[53, 54, 49], the MP2-based natural orbitals are the preferred choice for ground-state coupled cluster calculations due to their favorable computational cost.
The construction of natural spinors within the MP2 theory can be done as follows. The virtual-virtual block of 1-RDM $(\mathbf{D}^{\mathrm{(vv)}})$ in the MP2 method with matrix elements
$D_{ab}^{\mathrm{(vv)}}=-\sum_{c,i>j}\frac{\langle ac||ij\rangle{\langle bc||ij\rangle}}{\epsilon_{ij}^{ac}\epsilon_{ij}^{bc}}$ (33)
@@ II.6 Implementation and Computational details
The new relativistic qUCCSD[T] method has been implemented into the development version of our in-house quantum chemistry software package BAGH[55]. The package is written primarily in Python while the performance-critical components are optimized using Cython and Fortran. BAGH is interfaced with PYSCF[56, 57, 58], GAMESS US[59], socutils[60], and Dirac[61]. The X2CAMF-HF calculations are carried out using the socutils package[60], interfaced with BAGH. More details about the CD-X2CAMF-based implementation of the relativistic coupled cluster method can be found in ref[48]. Three predefined settings of the FNS and CD thresholds can be used: LOOSEFNS (FNS threshold: $10^{-4}$, and CD threshold: $10^{-3}$), NORMALFNS (FNS threshold: $10^{-4.5}$, and CD threshold: $10^{-4}$), and TIGHTFNS (FNS threshold: $10^{-5}$, and CD threshold: $10^{-5}$).The LOOSEFNS setting has been found to be appropriate for rapid estimation, normal FNS for standard energy difference, and the TIGHTFNS setting has been found to be appropriate for finite-field calculation of properties[48].
@@ III Result and Discussion
III.1 comparison with full CI and other triples correction schemes to unitary coupled cluster method in the non-relativistic regime
Bartlett and co-workers[62] have recently compared the various triples correction schemes to unitary coupled cluster methods with full CI results both for classical and quantum computers. As the present implementation of qUCCSD[T][T] is restricted to classical computers only, we have only chosen the variants that can be efficiently solved on classical computers. STO-6G basis set was used for the calculations together with the frozen core approximation following Ref.[62]. The UCC(3), UCC(4), UCCSD(4)$[T]$ and FCI results are taken from reference[62]. Figure 1 shows a schematic description of the error with respect to the Full CI results, and the corresponding values are presented in table 1. The errors are reported in milli Hartree(mH). CD and FNS approximations have not been used for the non-relativistic calculations. Although it can be seen that UCC(3), qUCCSD, and CCSD give very similar performance as compared to the FCI method, the performance of the partial triples correction scheme shows considerable deviation among each other. The UCC(4) shows the worst performance with an error as high as -31.46 mH. The UCCSD(4)$[T]$ method gives slightly better performance with a maximum error of 29.77 mH. The newly developed qUCCSD[T][T] method gives the best performance with a maximum error of - 7.728 mH, and its performance is comparable to the standard CCSD(T) method.
Figure 1: The error with respect to the FCI results in different approximate variants of unitary coupled cluster and standard coupled cluster for the total molecular energy.
Table 1: Error with respect to FCI in mH calculated in different method in STO-6G basis set and employing the frozen core approximation.
@@ III.2 Dissociation Enthalpy of Heavy metal-ligand complexes
To evaluate the practical applicability and accuracy of the qUCCSD[T] method in the relativistic domain, we computed the bond dissociation enthalpies (BDEs) of a representative set of 18 coinage metal-ligand complexes involving Cu+, Ag+, and Au+ ions. This test set was taken from our FNS-X2CAMF-CC implementation paper[48]. These systems are particularly challenging due to the combined influence of strong electron correlation and relativistic effects. All calculations were performed using the FNS-CD-X2CAMF framework, as detailed in Section IIF, and the NORMALFNS truncation threshold has been used for the calculation. The optimized geometries for all systems were taken from the study by Cavallo and co-workers [63]. For the ligand atoms, the aug-cc-pVXZ(X=D,T and Q) has been used. On the other hand, the dyall.aexz basis sets (x = 2, 3, and 4) were used for the metal cations Cu+, Ag+, and Au+. All basis sets were used in their uncontracted forms, and the frozen core approximation was used. The calculated HF and correlation energies were extrapolated using the three-point extrapolation by Peterson and Dunning.[64]
@@
Table 2: Experimental bond dissociation enthalpies for metal-ligand complexes (with error bars), along with those computed using the FNS-CD-X2CAMF based CCSD, qUCCSD, CCSD(T) and qUCCSD[T]] at complete basis set (CBS) limit.
Reactions CBS-CCSD[48] CBS-qUCCSD CBS-CCSD(T)[48] CBS-qUCCSD[T] Exp.[65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75]
@@
$\mathrm{(Cu\cdot C_{2}H_{6}O)}^{+}\rightarrow\mathrm{Cu}^{+}+\mathrm{C_{2}H_{6}O}$ 45.35 45.25 47.59 47.51 44.5 ± 2.9
@@
Figure 2: Experimental bond dissociation enthalpies (with error bars), along with those computed using the FNS-CD-X2CAMF-based CCSD, qUCCSD, CCSD(T), and qUCCSD[T]] method at complete basis set (CBS) limit.
@@ III.3 Bond Length and Vibrational Frequency
To further assess the performance of the relativistic qUCCSD[T] method, we benchmarked its ability to predict equilibrium bond lengths and harmonic vibrational frequencies for a series of diatomic molecules (HF, HCl, HBr, and HI), incorporating both light and heavy elements. The calculations were performed using uncontracted aug-cc-pVTZ (tz) and aug-cc-pVQZ (qz) basis sets for lighter atoms (H, F, Cl) and dyall.acv3z/dyall.acv4z for heavier atoms (Br, I).The bond length and harmonic vibrational frequency are calculated by numerical differentiation of the total energy using the TWOFIT utility program of DIRAC[61]. A fifth-order polynomial is used, and the TIGHTFNS threshold has been used for the calculations.
Table 3 presents the bond lengths obtained from CCSD, CCSD(T), qUCCSD, and qUCCSD[T], compared to experimental values[76]. The change from tz to qz level is small, so one could consider the qz results to be almost converged with respect to the basis set. The qUCCSD[T] method yields bond distances that are in excellent agreement with experiment and nearly indistinguishable from those predicted by CCSD(T), in addition to the advantage of its Hermitian formulation and consistent inclusion of spin–orbit effects via the X2CAMF Hamiltonian.
@@
. Molecule CCSD qUCCSD CCSD(T) qUCCSD[T] Exp.[76] tz qz tz qz tz qz tz qz HF 0.9180 0.915 0.9177 0.9149 0.9206 0.9118 0.9206 0.9178 0.9168 HCl 1.2752 1.2773 1.2760 1.2751 1.2776 1.2751 1.2776 1.2773 1.2745 HBr 1.4090 1.4095 1.4089 1.4094 1.4117 1.4132 1.4117 1.4122 1.4144 HI 1.6030 1.6033 1.6029 1.6031 1.6061 1.6068 1.6061 1.6066 1.6092
Table 4: Comparison of harmonic vibrational frequency (in $cm^{-1}$) calculated using the UCC-based method (qUCCSD and qUCCSD[T]) with the standard coupled cluster method (CCSD and CCSD(T)) under a fixed FNS and CD threshold $10^{-5}$. [^4] Uncontracted aug-cc-pVTZ (tz) and aug-cc-pVQZ (qz) basis sets were used for H, F, and Cl atoms, while dyall.acv3z (tz) and dyall.acv4z (qz) were employed for Br and I .
@@ III.4 Ionziation Potential and Electron Affinity
To assess the performance of the qUCCSD[T] method for electronic properties sensitive to both correlation and relativistic effects, we calculated vertical ionization potentials (IP) and electron affinities (EA) for a series of hydrogen halides (HF, HCl, HBr, HI, HAt) and group 13 heavy elements (In, Tl, Nh), respectively. All calculations were performed using uncontracted aug-cc-pVTZ and aug-cc-pVQZ basis sets for lighter elements and dyall.ae3z/ae4z basis sets for heavier elements. The s-aug-dyall.ae3z/ s-aug-dyall.ae4z basis set has been used for the heavy elements in the electron affinity calculations. The TIGHTFNS setting has been used for all the calculations. Table 5 shows that the qUCCSD[T] method yields ionization potentials that are highly consistent with experimental data[77, 78, 79, 80, 81], and often identical with CCSD(T) values. For example, the calculated IP of HCl using the qUCCSD[T] method is 12.71 eV (qz basis), very close to the experimental value of 12.74 ± 0.009 eV. Similar accuracy is observed for the heavier halides such as HBr and HI, confirming the efficacy of the relativistic framework. In almost all cases, the inclusion of partial triples correction improved upon the qUCCSD method, except for the HF molecule, where the results are of similar quality.
Table 6 presents electron affinities, where qUCCSD[T] again demonstrates good agreement with experimental references. For example, the EA of Tl calculated using qUCCSD[T] is 0.2828 eV (qz), compared to the experimental value of 0.32005 eV. While a slight underestimation is observed, the overall trend across In, Tl, and Nh is well-reproduced, highlighting the robustness of the qUCCSD[T] approach for systems with substantial spin–orbit coupling. The inclusion of the triples correction significantly improves over the qUCCSD results and gives a value comparable to the CCSD(T) method.
These results underline the capacity of the qUCCSD[T] method to deliver balanced accuracy even for energy differences, extending its applicability beyond thermochemistry and structural predictions.
Table 5: Comparison of ionization potential (IP) (in eV ) calculated using the UCC-based method (qUCCSD and qUCCSD[T]) with standard coupled cluster method (CCSD and CCSD(T)) under a fixed FNS and CD threshold $10^{-5}$. [^6] Uncontracted aug-cc-pVTZ (tz) and aug-cc-pVQZ (qz) basis sets were used for H, F, and Cl atoms, while dyall.acv3z (tz) and dyall.acv4z (qz) were employed for Br, I and At .
@@
HF 16.04 16.11 16.03 16.10 16.14 16.20 16.13 16.20 16.12±0.04[77]
HCl 12.55 12.64 12.55 12.64 12.62 12.72 12.62 12.71 12.74±0.009[79]
HBr 11.47 11.59 11.47 11.58 11.54 11.62 11.54 11.64 11.68±0.03[81]
HI 10.16 10.26 10.16 10.26 10.23 10.34 10.23 10.34 10.39±0.001[78]
HAt 9.06 9.18 9.05 9.17 9.12 9.24 9.12 9.24 9.317[80]
Table 6: Comparison of electron affinity (EA) (in eV ) calculated using the UCC-based method (qUCCSD and qUCCSD[T]) with the standard coupled cluster method (CCSDCCSD and $CCSD(T)$) under a fixed FNS and CD threshold $10^{-5}$.888s-aug-dyall.ae3z (tz) and s-aug-dyall.ae4z (qz) basis sets were used for the calculation.
@@
Nh 0.5351 0.6114 0.5367 0.6137 0.6040 0.6833 0.6043 0.6849 0.776  [^9] Theoretical best estimate (DCB +CCSDTQ+QED)[84]

### v2

@@ II.4 Cholesky Decomposition
In the CD framework,[45] the two-electron repulsion integrals (ERI) can be efficiently approximated as,
@@
where $\mu$, $\nu$, $\kappa$, $\lambda$ correspond to atomic spinor indices, $L_{\mu\nu}^{P}$ denotes the Cholesky vectors (CVs) and $n_{CD}$ specifies their dimensionality. Both single-step and double-step algorithms[46, 47, 48] have been proposed to perform CD of ERIs efficiently. In this present work, we employ the standard single-step approach, in which CVs are generated iteratively by selecting the largest diagonal elements of the ERI matrix $(\mu\nu\mid\mu\nu)$. The procedure is repeated until the largest diagonal elements fall below a predefined Cholesky threshold $\tau$, which determines the accuracy of the resulting decomposition.
The computed CVs on the AO basis can be transformed to the MO basis as follows,
@@
Using these transformed vectors, anti-symmetrized two electron integrals can be efficiently generated on the fly within the molecular orbital (MO) basis.
@@ II.5 Frozen Natural Spinors (FNS)
Relativistic CC calculations are more expensive than their non-relativistic counterparts due to the loss of spin symmetry and the additional need to store and process complex-valued quantities. Moreover, one often needs to uncontract the basis set to fit the small component of the Hamiltonian in relativistic calculations, which further increases the cost of relativistic CC calculations. One of the most effective ways to reduce the computational cost of relativistic CC calculations is to truncate the virtual space using natural spinors.[49, 50] The natural spinors are the relativistic analogs of the natural orbitals introduced by Löwdin[51] and can be obtained by diagonalizing the one-body correlated reduced density matrix (1-RDM), which is constructed from a spin-orbit coupled wave function calculated through some approximate electron correlation calculations.[52] Among the various flavors of natural spinors,[53, 54, 49] the MP2-based natural spinors are the preferred choice for ground-state coupled cluster correlated energy calculations due to their favorable computational cost.
The construction of natural spinors using the MP2 method can be performed as follows. The virtual-virtual block of 1-RDM $(\mathbf{D}^{\mathrm{(vv)}})$ in the MP2 method can be expressed as
$D_{ab}^{\mathrm{(vv)}}=-\frac{1}{2}\sum_{c,i>j}\frac{\langle ac||ij\rangle{\langle bc||ij\rangle}}{\epsilon_{ij}^{ac}\epsilon_{ij}^{bc}}$ (33)
@@ II.6 Implementation and Computational details
The new relativistic qUCCSD[T] method has been implemented into the development version of our in-house quantum chemistry software package BAGH.[55] The package is written primarily in Python while the performance-critical components are optimized using Cython and Fortran. BAGH is interfaced with PYSCF,[56, 57, 58] GAMESS US,[59] socutils,[60] and Dirac.[61] The X2CAMF-HF calculations are carried out using the socutils package,[60] interfaced with BAGH. More details about the CD-X2CAMF-based implementation of the relativistic coupled cluster method can be found in Ref. 62. Three predefined settings of the FNS and CD thresholds can be used: LOOSEFNS (FNS threshold: $10^{-4}$, and CD threshold: $10^{-3}$), NORMALFNS (FNS threshold: $10^{-4.5}$, and CD threshold: $10^{-4}$), and TIGHTFNS (FNS threshold: $10^{-5}$, and CD threshold: $10^{-5}$).The LOOSEFNS setting has been found to be appropriate for rapid estimation, normal FNS for standard energy difference, and the TIGHTFNS setting has been found to be appropriate for finite-field calculation of properties.[62]
@@ III Result and Discussion
III.1 comparison with full CI and other triples correction schemes to UCC method in the non-relativistic regime
Bartlett and co-workers[17] have recently compared the various triples correction schemes to UCC methods with FCI results both for classical and quantum computers. As the present implementation of qUCCSD[T][T] is restricted to classical computers only, we have chosen to compare with the variants that can be efficiently solved on classical computers. STO-6G basis set was used for the calculations together with the frozen core approximation following Ref. 17. The UCC(3), UCC(4), UCCSD(4)$[T]$, CCSD, CCSD(T) and FCI results are taken from Ref. 17. Figure 1 shows a schematic description of the error with respect to the FCI results, and the corresponding values are presented in table 1. The errors are reported in milli Hartree(mH). CD and FNS approximations have not been used for the non-relativistic calculations. Although it can be seen that UCC(3), qUCCSD, and CCSD give very similar performance as compared to the FCI method, the performance of the partial triples correction scheme shows considerable deviation among each other. The UCC(4) shows the worst performance with an error as high as -36.78 mH. The UCCSD(4)$[T]$ method gives slightly better performance with a maximum error of 29.77 mH. The newly developed qUCCSD[T][T] method gives the best performance with a maximum error of - 7.728 mH, and its performance is comparable to the standard CCSD(T) method.
Figure 1: Deviation in total molecular energies (in mH) from FCI results using the frozen core approximation for various approximate unitary and standard coupled cluster methods.
Table 1: Errors (mH) with respect to FCI calculated using various methods in STO-6G basis set with the frozen core approximation.
@@ III.2 Dissociation Enthalpy of Heavy metal-ligand complexes
To evaluate the practical applicability and accuracy of the qUCCSD[T] method in the relativistic domain, we computed the bond dissociation enthalpies (BDEs) of a representative set of 18 coinage metal-ligand complexes involving Cu+, Ag+, and Au+ ions. This test set was taken from Ref. 62. These systems are particularly challenging due to the combined influence of strong electron correlation and relativistic effects. All calculations were performed using the FNS-CD-X2CAMF framework, as detailed in Section IIF, and the NORMALFNS truncation threshold has been used for the calculation. The optimized geometries for all systems were taken from the study by Cavallo and co-workers.[63] For the ligand atoms, the aug-cc-pVXZ(X=D,T and Q) has been used. On the other hand, the dyall.aexz basis sets (x = 2, 3, and 4) were used for the metal cations Cu+, Ag+, and Au+. All basis sets were used in their uncontracted forms, and the frozen core approximation was used. The calculated HF and correlation energies were extrapolated using the three-point extrapolation by Peterson and Dunning.[64]
@@
Table 2: Experimental bond dissociation enthalpies (kcal/mol) for metal-ligand complexes (with error bars), along with those computed using the FNS-CD-X2CAMF based CCSD, qUCCSD, CCSD(T) and qUCCSD[T]] at complete basis set (CBS) limit.
Reactions CBS-CCSD[62] CBS-qUCCSD CBS-CCSD(T)[62] CBS-qUCCSD[T] Exp.[65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76]
@@
$\mathrm{(Cu\cdot C_{2}H_{6}O)}^{+}\rightarrow\mathrm{Cu}^{+}+\mathrm{C_{2}H_{6}O}$ 45.35 45.27 47.59 48.23 44.5 ± 2.9
@@
Figure 2: Experimental bond dissociation enthalpies (with error bars), along with those computed using the FNS-CD-X2CAMF-based CCSD, qUCCSD, CCSD(T), and qUCCSD[T] method at complete basis set (CBS) limit.
@@ III.3 Bond Length and Vibrational Frequency
To further assess the performance of the relativistic qUCCSD[T] method, we benchmarked its ability to predict equilibrium bond lengths and harmonic vibrational frequencies for a series of diatomic molecules (HF, HCl, HBr, and HI), incorporating both light and heavy elements. The calculations were performed using uncontracted aug-cc-pVTZ (tz) and aug-cc-pVQZ (qz) basis sets for lighter atoms (H, F, Cl) and dyall.acv3z/dyall.acv4z for heavier atoms (Br, I).The bond length and harmonic vibrational frequency are calculated by numerical differentiation of the total energy using the TWOFIT utility program of DIRAC.[61] A fifth-order polynomial is used, and the TIGHTFNS threshold has been used for the calculations.
Table 3 presents the bond lengths obtained from CCSD, CCSD(T), qUCCSD, and qUCCSD[T], compared to experimental values.[77] The change from tz to qz level is small, so one could consider the qz results to be almost converged with respect to the basis set. The qUCCSD[T] method yields bond distances that are in excellent agreement with experiment and nearly indistinguishable from those predicted by CCSD(T), in addition to the advantage of its Hermitian formulation and consistent inclusion of spin–orbit effects via the X2CAMF Hamiltonian.
@@
. Molecule CCSD qUCCSD CCSD(T) qUCCSD[T] Exp.[77] tz qz tz qz tz qz tz qz HF 0.9178 0.9151 0.9177 0.9149 0.9207 0.9179 0.9205 0.9178 0.9168 HCl 1.2760 1.2751 1.2760 1.2751 1.2782 1.2774 1.2781 1.2773 1.2745 HBr 1.4090 1.4095 1.4089 1.4095 1.4117 1.4123 1.4116 1.4122 1.4144 HI 1.6030 1.6033 1.6029 1.6031 1.6061 1.6067 1.6060 1.6066 1.6092
Table 4: Comparison of harmonic vibrational frequency ($cm^{-1}$) calculated using the UCC-based method (qUCCSD and qUCCSD[T]) with the standard coupled cluster method (CCSD and CCSD(T)) under a fixed FNS and CD threshold $10^{-5}$ with frozen core approximation.
@@ III.4 Ionziation Potential and Electron Affinity
To assess the performance of the qUCCSD[T] method for electronic properties sensitive to both correlation and relativistic effects, we calculated vertical ionization potentials (IP) and electron affinities (EA) for a series of hydrogen halides (HF, HCl, HBr, HI, HAt) and group 13 heavy elements (In, Tl, Nh), respectively. All calculations were performed using uncontracted aug-cc-pVTZ and aug-cc-pVQZ basis sets for lighter elements and dyall.ae3z/ae4z basis sets for heavier elements. The s-aug-dyall.ae3z/ s-aug-dyall.ae4z basis set has been used for the heavy elements in the EA calculations. The TIGHTFNS setting has been used for all the calculations. Table 5 shows that the qUCCSD[T] method yields IPs that are highly consistent with experimental data,[78, 79, 80, 81, 82] and often identical with CCSD(T) values. For example, the calculated IP of HCl using the qUCCSD[T] method is 12.718 eV (qz basis), very close to the experimental value of 12.74 ± 0.009 eV. Similar accuracy is observed for the heavier halides such as HBr and HI, confirming the efficacy of the relativistic framework. In almost all cases, the inclusion of partial triples correction improved upon the qUCCSD method, except for the HF molecule, where the results are of similar quality.
Table 6 presents EAs, where qUCCSD[T] again demonstrates good agreement with experimental references. For example, the EA of Tl calculated using qUCCSD[T] is 0.2828 eV (qz), compared to the experimental value of 0.32005 eV. While a slight underestimation is observed, the overall trend across In, Tl, and Nh is well-reproduced, highlighting the robustness of the qUCCSD[T] approach for systems with substantial spin–orbit coupling. The inclusion of the triples correction significantly improves over the qUCCSD results and gives a value comparable to the CCSD(T) method.These results underline the capacity of the qUCCSD[T] method to deliver balanced accuracy even for energy differences, extending its applicability beyond thermochemistry and structural predictions.
Table 5: Comparison of ionization potential (eV ) calculated using the UCC-based method (qUCCSD and qUCCSD[T]) with standard $C$-based method (CCSD and CCSD(T)) under a fixed FNS and CD threshold $10^{-5}$ with frozen core approximation.
@@
HF 16.043 16.106 16.038 16.101 16.137 16.200 16.135 16.198 16.12±0.04[78]
HCl 12.551 12.639 12.552 12.640 12.625 12.719 12.624 12.718 12.74±0.009[80]
HBr 11.473 11.581 11.473 11.580 11.542 11.643 11.541 11.642 11.68±0.03[82]
HI 10.162 10.260 10.163 10.260 10.229 10.337 10.228 10.337 10.39±0.001[79]
HAt 9.054 9.175 9.053 9.173 9.121 9.240 9.121 9.239 9.317[81]
Table 6: Comparison of electron affinity (eV ) obtained using the UCC-based methods (qUCCSD and qUCCSD[T]) and the standard $C$-based methods (CCSDCCSD and $CCSD(T)$), with a fixed FNS and CD threshold of $10^{-5}$ and $10^{-4.5}$ respectively. s-aug-dyall.ae3z (tz) and s-aug-dyall.ae4z (qz) basis sets were used for the calculations.
@@
Nh 0.5351 0.6114 0.5367 0.6137 0.6040 0.6849 0.6043 0.6849 0.776  [^1] Theoretical best estimate (DC(B) +CCSDTQ+QED)[85]